### v1

# Towards bilipschitz extension from euclidean separated nets of general dimension

Abstract
We provide a foundation for an approach to the open problem of bilipschitz extendability of mappings defined on a Euclidean separated net. In particular, this allows for the complete positive solution of the problem in dimension two. Along the way, we develop a set of tools for bilipschitz extensions of mappings between subsets of Euclidean spaces.
1 Introduction
The purpose of the present article is to lay foundations for an approach to the problem of whether every bilipschitz mapping of a separated net of $\mathbb{R}^{d}$ to $\mathbb{R}^{d}$ admits a bilipschitz extension to the whole of $\mathbb{R}^{d}$: a 23 year old open question (in all dimensions $d\geq 2$) posed by Alestalo, Trotsenko and Väisälä in [1]. Indeed the framework provided in the present work allows for the complete positive solution of this bilipschitz extension problem in dimension 2: In a companion article [10] we perform in detail various two-dimensional constructions which, together with the hereafter established results, provide for any given $L$-bilipschitz mapping of a separated net of $\mathbb{R}^{2}$ to $\mathbb{R}^{2}$ a $C(L)$-bilipschitz extension to the whole of $\mathbb{R}^{2}$. [^1] We have decided to split [10] into two papers; the present paper contains all its general dimensional results with improved statements and significantly simplified and streamlined proofs. A new resubmision of [10] will contain only the 2-dimensional arguments needed to finish the proof of the bilipschitz extendability of any bilipschitz mapping on $\mathbb{Z}^{2}$.
There is a pantheon of mathematical research going back to the early 20th century on the general problem of extending a given mapping of a subset of an ambient space whilst preserving its key properties. We highlight the famous extension theorems of Tietze–Urysohn [30] for continuous and Whitney [31] for smooth functions. The class of Lipschitz functions/mappings lies between the continuous and smooth classes. Due primarily to Rademacher’s theorem [23], the Lipschitz condition suffices in Euclidean settings to prevent the pathological behaviour permissible under just continuity. On the other hand the Lipschitz condition allows for much greater flexibility in comparison with smooth: note that important natural mappings such as the Euclidean norm (and many other norms and related functions such as distance functions and projections) are Lipschitz but not smooth. Unlike both continuity or differentiability, the Lipschitz condition transfers seamlessly to discrete settings as well as to general metric spaces. Further, Lipschitz mappings play an important role in the theory of discrete metric spaces, for example in the study of embeddings [19, 22] or in the notion of bilipschitz equivalence, e.g., [7, 20].
Lipschitz mappings are well extendable in some key settings. Kirszbraun’s theorem [12] states that any Lipschitz mapping of a subset of Hilbert space to another Hilbert space may extended to the whole space, even without increasing its Lipschitz constant. Moreover, McShane’s Extension Theorem [21] similarly provides Lipschitz extensions with the same Lipschitz constant for any (real-valued) Lipschitz function of a subset of a metric space. Kirszbraun’s theorem continues to shape the modern field; we point to two significant improvements [14] and [5].
That the analogue of Kirszbraun’s theorem for bilipschitz mappings fails is, after some thought, obvious, and the reason is topological. For example we may consider a bilipschitz mapping of the unit sphere and the origin, which maps the origin ‘outside’ the image of the sphere. Such a mapping cannot even be extended to a homeomorphism of the space, let alone to a bilipschitz one. Thus research on the bilipschitz extension problem has focussed on the question of when we may extend a given bilipschitz mapping of a subset. From what kind of subsets is this possible? Most existing positive results occur in dimension two; see [26, 27, 8, 15, 16] which, in particular, provide bilipschitz extensions from a line and from the boundary of a square (or circle). We will discuss the two dimensional situation in more detail in the companion article [10].
The challenge of bilipschitz extension in higher dimensional spaces is made evident by the number of pathological examples and notorious enduring open problems connected to extendability of mappings between higher dimensional spaces. The Jordan–Schoenflies Theorem (see, e.g., [25, Thm. 3.1]) states that any embedding of $S^{1}$ into $\mathbb{R}^{2}$ may be extended to a homeomorphism of $\mathbb{R}^{2}$. However, this statement fails for embeddings of $S^{d-1}$ into $\mathbb{R}^{d}$ in all higher dimensions $d\geq 3$: in dimension 3 a counterexample is provided by Alexander’s Horned Sphere [2]. Shockingly the statement still fails if the given embedding is also assumed to be bilipschitz. A counterexample is given by the Fox-Artin Wild Sphere [18, Theorem 3.7].
Using the fact that the sphere $S^{d-1}$ minus a point is homeomorphic to $\mathbb{R}^{d-1}$, the Jordan–Schoenflies theorem can be reformulated for embeddings of $\mathbb{R}$ into $\mathbb{R}^{2}$ instead of $S^{1}$ into $\mathbb{R}^{2}$. Further, when the given embedding of $\mathbb{R}$ is $L$-bilipschitz, the extension to $\mathbb{R}^{2}$ can be made $C(L)$-bilipschitz [26, 15]. This evolution of the Jordan–Schoenflies Theorem has particular relevance to the problem of extending a given bilipschitz mapping $f\colon X\to\mathbb{R}^{d}$ of a separated net $X$ of $\mathbb{R}^{d}$.
[…]
To overcome the obstacles remaining in the higher dimensional setting will require significant future research. The present article aims to be a start along this path. We now state our main results and explain how they can form part of a plausible programme to solve the higher dimensional problem.
Main Results
Alestalo, Trotsenko and Väisälä [1, 4.4(ii)] pose the questions, in dimensions $d\geq 2$, of the bilipschitz extendability of a given bilipschitz mapping $f$ of
  1. (a)
    $\mathbb{Z}^{d}$ to $\mathbb{R}^{d}$,
  2. (b)
    a given separated net of $\mathbb{R}^{d}$ to $\mathbb{R}^{d}$.
Whilst (a) is clearly a subquestion of (b), [1] points out that any further relationship between these questions is unknown and far from obvious. The naive argument to try to show that questions (a) and (b) are equivalent would require, for a given separated net $X\subseteq\mathbb{R}^{d}$, there to be a bilipschitz bijection between $X$ and $\mathbb{Z}^{d}$. However, there exist separated nets not admitting any such bijection; these are called ‘non-rectifiable’ or ‘not bilipschitz equivalent to the lattice’ in the literature; see [7, 20]. Despite this occurrence of ‘very different’ separated nets of $\mathbb{R}^{d}$, our first main result reduces, in all dimensions, the bilipschitz extension problem for all separated nets to that of the integer lattice, establishing the equivalence of questions (a) and (b) of Alestalo, Trotsenko and Väisälä [1, 4.4(ii)].
Theorem 1.1.
Let $d\in\mathbb{N}$. Then the following are equivalent:
  1. (i)
    Every bilipschitz mapping $f\colon\mathbb{Z}^{d}\to\mathbb{R}^{d}$ admits a bilipschitz mapping $F\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ with $F|_{\mathbb{Z}^{d}}=f$.
  2. (ii)
    For every separated net $A$ of $\mathbb{R}^{d}$ and every bilipschitz mapping $f\colon A\to\mathbb{R}^{d}$ there is a bilipschitz mapping $F\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ with $F|_{A}=f$.
Furthermore, if (i) holds with $\operatorname{bilip}(F)\leq C\left(d,\operatorname{bilip}(f)\right)$ for some monotone increasing function $C_{d}\colon[1,\infty)\to[1,\infty)$ then (ii) holds with
$\operatorname{bilip}(F)\leq K\cdot C_{d}\left(24\sqrt{d}R^{2}K^{3}\operatorname{bilip}(f)\right),$
where $K:=16\max\left\{\frac{3d}{r},1\right\}$ and $R,r$ denote the net and separation constants of $A$ respectively.
Our second main result characterises the bilipschitz extension problem on the integer lattice in terms of existence of an ‘intermediate extension’ to a sequence of ‘horizontal’ hyperplanes.
Theorem 1.2.
Let $d\in\mathbb{N}$, $d\geq 2$ and $f\colon\mathbb{Z}^{d}\to\mathbb{R}^{d}$ be a bilipschitz mapping. Then the following are equivalent:
  1. (i)
    There exists a bilipschitz mapping $F\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $F|_{\mathbb{Z}^{d}}=f$.
  2. (ii)
    There exist $T\in\mathbb{N}$ and a bilipschitz mapping $G\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that the mapping
    $\widetilde{F}\colon\mathbb{Z}^{d}\cup\left(\mathbb{R}^{d-1}\times\left(T\mathbb{Z}+\frac{1}{2}\right)\right)\to\mathbb{R}^{d},\qquad\widetilde{F}(x)=\begin{cases}f(x)&\text{if }x\in\mathbb{Z}^{d},\\ G(x)&\text{if }x\in\mathbb{R}^{d-1}\times\left(T\mathbb{Z}+\frac{1}{2}\right),\end{cases}$
    is bilipschitz and for every $k\in\mathbb{Z}$ it holds that
    $f\left(\mathbb{Z}^{d}\cap\left(\mathbb{R}^{d-1}\times\left[(k-1)T+\frac{1}{2},kT+\frac{1}{2}\right]\right)\right)\subseteq G\left(\mathbb{R}^{d-1}\times\left[(k-1)T+\frac{1}{2},kT+\frac{1}{2}\right]\right).$
Furthermore, there is a monotone increasing function $\beta_{d}\colon[1,\infty)\to[1,\infty)$ with $2\lfloor t\rfloor-1\leq\beta_{d}(t)\leq\exp(8dt)$ for all $t\in[1,\infty)$ such that whenever (ii) holds with $M_{1}\geq\operatorname{bilip}(\widetilde{F})$ and $M_{2}\geq\operatorname{bilip}(G)$, the bilipschitz extension $F$ of $f$ in (i) may be found with
$\operatorname{bilip}(F)\leq\beta_{d}\left(2^{42}d^{9/2}\left(M_{1}M_{2}\right)^{10}T^{3+\frac{3}{d-1}}\right)^{3^{d-1}+1}.$
Moreover, the function $\beta_{d}$ may be defined as in Lemma 6.
Future Programme
Now we explain how Theorem 1.2 can become an important stepping stone in proving that every bilipschitz mapping $f\colon\mathbb{Z}^{d}\to\mathbb{R}^{d}$ admits a bilipschitz extension $\mathbb{R}^{d}\to\mathbb{R}^{d}$. We describe a strategy to obtain the mappings $\widetilde{F}$ and $G$ from the statement (ii) of Theorem 1.2 in general dimension $d\geq 2$; we emphasise that while we complete all of the proposed steps in the case $d=2$ in [10], for $d\geq 3$ some of the steps remain currently open.
For convenience, we denote the hyperplane $\mathbb{R}^{d-1}\times\left\{h\right\}$, $h\in\mathbb{R}$, by $\mathcal{H}_{h}$. Assume we are given a bilipschitz mapping $f\colon\mathbb{Z}^{d}\to\mathbb{R}^{d}$. We suggest the following steps:
  1. (1)
    Construct a bilipschitz extension $F_{0}\colon\mathbb{Z}^{d}\cup\mathcal{H}_{\frac{1}{2}}\to\mathbb{R}^{d}$ in such a way that there is a bilipschitz mapping $G_{0}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ so that $G_{0}|_{\mathcal{H}_{\frac{1}{2}}}=F_{0}|_{\mathcal{H}_{\frac{1}{2}}}$.
  2. (2)
    Adjust the extension $F_{0}$ so that, in addition to its property from (1), it separates the points of $\mathbb{Z}^{d}$ correctly, i.e., $x,y\in\mathbb{Z}^{d}$ belong to the same connected component of $\mathbb{R}^{d}\setminus\mathcal{H}_{\frac{1}{2}}$ if and only if $F_{0}(x),F_{0}(y)$ belong to the same connected component of $\mathbb{R}^{d}\setminus F_{0}\left(\mathcal{H}_{\frac{1}{2}}\right)$.
  3. (3)
    For a parameter $T=T(\operatorname{bilip}(f))$, construct a bilipschitz mapping
    $\widetilde{F}\colon\mathbb{Z}^{d}\cup\bigcup_{z\in\mathbb{Z}}\mathcal{H}_{\frac{1}{2}+Tz}\to\mathbb{R}^{d}$
    extending $f$ such that for every $z\in\mathbb{Z}$ both (1) and (2) hold for each $\mathcal{H}_{\frac{1}{2}+Tz}$ separately: that is, for every $z\in\mathbb{Z}$ there is a bilipschitz mapping $G_{z}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ so that $G_{z}|_{\mathcal{H}_{\frac{1}{2}+Tz}}=\widetilde{F}|_{\mathcal{H}_{\frac{1}{2}+Tz}}$, and moreover, $x,y\in\mathbb{Z}^{d}$ belong to the same connected component of $\mathbb{R}^{d}\setminus\mathcal{H}_{\frac{1}{2}+Tz}$ if and only if $\widetilde{F}(x),\widetilde{F}(y)$ belong to the same connected component of $\mathbb{R}^{d}\setminus\widetilde{F}\left(\mathcal{H}_{\frac{1}{2}+Tz}\right)$.
  4. (4)
    Use the mapping $\widetilde{F}$ from (3) to produce a bilipschitz mapping $G\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ so that for every $z\in\mathbb{Z}$ it holds that $G|_{\mathcal{H}_{\frac{1}{2}+Tz}}=\widetilde{F}|_{\mathcal{H}_{\frac{1}{2}+Tz}}$.
It is not difficult to show that after Step (4) the mappings $\widetilde{F}$ and $G$ satisfy the assumptions of Theorem 1.2(ii).
In Step (2) we can make use of the easy observation that, following Step (1), there will be a parameter $P=P(\operatorname{bilip}(F_{0}))>0$ such that whenever $x,y\in\mathbb{Z}^{d}$ belong to the same connected component of $\mathbb{R}^{d}\setminus\mathcal{H}_{\frac{1}{2}}$ and $F_{0}(x),F_{0}(y)$ belong to different connected components of $\mathbb{R}^{d}\setminus F_{0}\left(\mathcal{H}_{\frac{1}{2}}\right)$ (or vice versa), then both $x,y$ are at distance at most $P$ from $\mathcal{H}_{\frac{1}{2}}$. Thus, transforming the picture by applying $G_{0}^{-1}$ from Step (1), we get that the image points on the ‘wrong side’ of $G_{0}^{-1}\circ F_{0}(\mathcal{H}_{\frac{1}{2}})=\mathcal{H}_{\frac{1}{2}}$ all occur within a horizontal slab around $\mathcal{H}_{\frac{1}{2}}$ of thickness determined by $\operatorname{bilip}(G_{0}^{-1}\circ F_{0})$. The modification of Step (2) can therefore take place within this horizontal slab; see [10, Theorem 7.1] for the 222-dimensional analogue of this. We believe that our 2-dimensional construction should also work in general dimension; however, there are significant technical difficulties in describing the relevant local reparametrizations of multi-dimensional mappings formally.
Step (3) is not difficult to implement provided one already knows how to perform both Steps (1) and (2); in these cases, the relevant considerations in [10, Lem. 9.3] generalise to $d$ dimensions straightforwardly. If the constructions for Steps (1) and (2) are performed separately for multiple hyperplanes, then taking these hyperplanes sufficiently far apart ensures that the separate constructions don’t meaningfully interact in the calculation of the global bilipschitz constant.
Steps (1) and (4) are completely open. The theory of bilipschitz extensions in general dimension is currently developed for mappings defined on spheres [24, 28, 17], but not for those defined on hyperplanes. Although the two settings are often very closely related through the stereographic projection, the latter is only locally bilipschitz, and thus, inconsequential in a bilipschitz setting. However, there are promising indications from the positive results available for the spherical case and one can follow the (highly sophisticated) proofs and try to obtain the analogous results for hyperplanes.
[…]
Speculatively assuming that the analogue of the bilipschitz annulus theorem holds true also when spheres are replaced by hyperplanes and annuli by slabs between two hyperplanes, we can use it after Step (3) to obtain a bilipschitz mapping $\Phi\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ so that $\Phi\circ\widetilde{F}\left(\mathcal{H}_{\frac{1}{2}+Tz}\right)=\mathcal{H}_{\frac{1}{2}+Tz}$ for every $z\in\mathbb{Z}$; the existence of the mappings $\left(G_{z}\right)_{z\in\mathbb{Z}}$ from Step (3) ensures that the local bilipschitz flattness condition from the annulus theorem is satisfied. Assuming further that also an analogue of the corollary [29, Thm. 3.14] mentioned above holds true with the sphere replaced by a hyperplane and the ball by a half-space, the sequence of bilipschitz mappings $\left(\left(\Phi\circ\widetilde{F}\right)|_{\mathcal{H}_{\frac{1}{2}+Tz}}\right)_{z\in\mathbb{Z}}$ can be simultaneously extended by a single bilipschitz mapping $\widetilde{G}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$. Then $\Phi^{-1}\circ\widetilde{G}$ gives the mapping $G$ sought after in Step (4).
Our strategy from [10, Sec. 4] to complete Step (1) for $d=2$ seems to be very specific to dimension 222. Unfortunately, we currently do not know how to extend it to higher dimensions. We only remark that, again in an analogy with the spherical setting, a sufficient condition ensuring the existence of the mapping $G_{0}$ from Step (1) should be the local bilipschitz flattness of $F_{0}|_{\mathcal{H}_{\frac{1}{2}}}$; we refer to the resolution of the so-called Schoenflies problem in the bilipschitz category in all dimensions $d$ from [17, Thm. 7.8].
Strategy for the proofs of the main theorems
Returning to the main results of the present paper, we describe the big picture of our proofs of Theorems 1.1 and 1.2, picking out the key tools and stepping stones established in the sections that follow.
Ultimately, every bilipschitz extension constructed in the present paper will be obtained by gluing and composing very simple bilipschitz transformations. These ‘building blocks’ will be bilipschitz mappings designed to perform a swap of two specified points, denoted $x$ and $y$ below, whilst disturbing the rest of the space as little as possible. The next lemma, proved in Section 3, provides these mappings with an estimate on their bilipschitz constant: {restatable*}lemmarestatespin Let $d\in\mathbb{N}$, $d\geq 2$, $x,y\in\mathbb{R}^{d}$ and $0<r\leq\frac{\left\|y-x\right\|}{2}$. Then there exists a $\frac{4\left\|y-x\right\|^{2}}{r^{2}}$-bilipschitz mapping $\tau\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that
  1. (i)
    $\tau(z)=z$ for all $z\in\mathbb{R}^{d}\setminus B([x,y],r)$.
  2. (ii)
    $\tau(y)=x$ and $\tau(x)=y$.
[…]
Proof of Theorem 1.1 (Sketch).
Theorem 1.1 is consequence of two facts:
  1. (a)
    Any $(r,R)$-separated net $A\subseteq\mathbb{R}^{d}$ may be transformed via a $P(d,r,R)$-bilipschitz mapping $\Phi\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ so that $\Phi(A)\subseteq\mathbb{Z}^{d}$.
  2. (b)
    For any $R$-net $Y\subseteq\mathbb{Z}^{d}$ and any $L$-bilipschitz mapping $f\colon Y\to\mathbb{R}^{d}$ there is a $P^{\prime}(d,R,L)$-bilipschitz extension of $f$ to $\mathbb{Z}^{d}$.
To verify the only non-trivial implication, (ii)$\Rightarrow$(i), of Theorem 1.1, let $f\colon A\to\mathbb{R}^{d}$ be given by (ii), $\Phi$ be given by (a) and then let (b) provide the bilipschitz extension $g\colon\mathbb{Z}^{d}\to\mathbb{R}^{d}$ of $f\circ\Phi^{-1}\colon Y\to\mathbb{R}^{2}$ with $Y=\Phi(A)\subseteq\mathbb{Z}^{2}$. The desired extension of $f\colon A\to\mathbb{R}^{d}$ is then given by $G\circ\Phi$ where $G$ is the bilipschitz extension of $g$ given by (i).
The proof of (a) is based on Lemmas 1 and 1. First we blow up the separated net $A$, so that its separation becomes large in comparison to $\sqrt{d}$. This ensures that every point of the transformed net $A^{\prime}$ has a unique choice of nearest point in $\mathbb{Z}^{d}$ and that these pairs of neighbours are well separated from each other. We then perform simultaneous swappings of the points of $A^{\prime}$ and their chosen nearest neighbours in $\mathbb{Z}^{d}$, by gluing together mappings $\tau$ from Lemma 1 using Lemma 1.
For (b) we calculate a radius $s$ so that the balls of radius $s$ around each point of $f(Y)$ are pairwise disjoint with some separation. Next we observe that the collection of balls $(B(y,R))_{y\in Y}$ cover $\mathbb{R}^{d}$ and we can estimate the total number of integer lattice points in each such ball. It remains to calculate a large enough integer $N$ so that for any $y,y^{\prime}\in Y$ the set $\left(f(y)+\frac{1}{N}\mathbb{Z}^{d}\right)\cap B(f(y),s)$ contains more points than in $\mathbb{Z}^{d}\cap B(y^{\prime},R)$. This allows for an injective mapping from $\mathbb{Z}^{d}\setminus Y$ to $\mathbb{R}^{d}\setminus f(Y)$ with the additional property that any $x$ in the domain admits $y\in Y$ with $x\in B(y,R)$ and $f(x)\in\left(f(y)+\frac{1}{N}\mathbb{Z}^{d}\right)\cap B(f(y),s)$. An arbitrary choice of such mapping determines a suitable extension of $f$. ∎
The proof of Theorem 1.2 will be based on the following theorem, proved in Section 6: {restatable*}thmrestatethread Let $d\in\mathbb{N}$ with $d\geq 2$. Then there is a monotone increasing function $\beta_{d}\colon[1,\infty)\to[1,\infty)$, with $2\lfloor t\rfloor-1\leq\beta_{d}(t)\leq\exp(8dt)$ for all $t\in[1,\infty)$, such that the following statement holds: For any $H\in\mathbb{N}$, $L\geq 1$ and any $L$-bilipschitz mapping
$f\colon\left(\mathbb{R}^{d-1}\times\left\{\frac{1}{2},H+\frac{1}{2}\right\}\right)\cup\left(\mathbb{Z}^{d-1}\times\left\{1,\ldots,H\right\}\right)\to\mathbb{R}^{d}$
satisfying
$f(x)=x\qquad\text{for all }x\in\mathbb{R}^{d-1}\times\left\{\frac{1}{2},H+\frac{1}{2}\right\}$
and
$f\left(\mathbb{Z}^{d-1}\times\left\{1,\ldots,H\right\}\right)\subseteq\mathbb{R}^{d-1}\times\left[\frac{1}{2},H+\frac{1}{2}\right],$
there is a bilipschitz extension $F\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ of $f$ with
$\operatorname{bilip}(F)\leq 2^{38}d^{9/2}L^{9}H^{2+\frac{3}{d-1}}\cdot\beta_{d}\left(2^{42}d^{9/2}L^{10}H^{3+\frac{3}{d-1}}\right)^{3^{d-1}}.$
Moreover, the function $\beta_{d}\colon[1,\infty)\to[1,\infty)$ may be defined as in Lemma 6.
Proof.
(Sketch) We aim to construct a bilipschitz mapping $\Theta\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that
  1. (i)
    $\Theta(x)=x$ for all $x\in\mathbb{R}^{d-1}\times\left(\mathbb{R}\setminus\left(\frac{1}{2},H+\frac{1}{2}\right)\right)$, and
  2. (ii)
    $\Theta\circ f(x,m)=(x,m)$ for all $m\in[H]$ and $x\in\mathbb{Z}^{d-1}$.
The bilipschitz $F\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ verifying Theorem 1 may then be defined by $F:=\Theta^{-1}$. The required transformation $\Theta$ will be built by composing and gluing together mappings of the form of $\tau$ from 1. By composing and gluing together these ‘swapping transformations’ $\tau$ we first construct a bilipschitz mapping $\Psi$ which approximately projects each ‘image layer’ $f(\mathbb{Z}^{d-1}\times\left\{m\right\})$ into a finer rescaling of the lattice $\mathbb{Z}^{d-1}\times\left\{m\right\}$, more specifically $\frac{1}{N}\mathbb{Z}^{d-1}\times\left\{m\right\}$ for some quantity $N\in\mathbb{N}$ controlled by $d$, $L$ and $H$. This is achieved by Lemma 7.
[…]
With Theorem 1 and Lemma 1 we can already give the full proof of Theorem 1.2:
Proof of Theorem 1.2.
It is clear that (i) implies (ii).
Let us now suppose that $T$, $G$ and $\widetilde{F}$ are given by (ii) and suppose $M_{1}\geq\operatorname{bilip}(\widetilde{F})$ $M_{2}\geq\operatorname{bilip}(G)$. For each $k\in\mathbb{Z}$ we set $A_{k}:=\mathbb{R}^{d-1}\times\left((k-1)T+\frac{1}{2},kT+\frac{1}{2}\right)$ and apply Theorem 1 (appropriately shifted) to the $M_{1}M_{2}$-bilipschitz mapping $G^{-1}\circ\widetilde{F}$ restricted to the set
$V_{k}:=\partial A_{k}\cup\left(\mathbb{Z}^{d-1}\times\left\{(k-1)T+1,(k-1)T+2,\ldots,kT\right\}\right).$
We obtain for each $k\in\mathbb{Z}$ a bilipschitz mapping $F_{k}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that
$F_{k}(x)=G^{-1}\circ\widetilde{F}(x)\quad\text{for all }x\in V_{k},\\ \text{and}\quad\operatorname{bilip}(F_{k})\leq 2^{38}d^{9/2}\left(M_{1}M_{2}\right)^{9}T^{2+\frac{3}{d-1}}\cdot\beta_{d}\left(2^{42}d^{9/2}\left(M_{1}M_{2}\right)^{10}T^{3+\frac{3}{d-1}}\right)^{3^{d-1}}=:C,$
where $\beta_{d}\colon[1,\infty)\to[1,\infty)$ may be defined by Lemma 6. We define a mapping $\Phi\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ as
$\Phi(x):=F_{k}(x)\qquad\text{ whenever }x\in\mathbb{R}^{d-1}\times\left[(k-1)T+\frac{1}{2},kT+\frac{1}{2}\right].$
The mapping $\Phi$ is well-defined, since $F_{k}|_{\mathbb{R}^{d-1}\times\left\{(k-1)T+\frac{1}{2},kT+\frac{1}{2}\right\}}$ is equal to the identity for every $k\in\mathbb{Z}$.
Applying Lemma 1 to $\Phi$, the collection of open sets $\left(A_{k}\right)_{k\in\mathbb{Z}}$ and $C$ defined above, we establish that $\Phi$ is $C$-bilipschitz. To complete the proof, it only remains to set $F:=G\circ\Phi$. To refine the upper bound $\operatorname{bilip}(F)\leq C\operatorname{bilip}(G)\leq CM_{2}$ to the form of the ‘Furthermore’ statement, we apply the monotonicity of $\beta_{d}$ and the inequalities $\beta_{d}(t)\geq 2\lfloor t\rfloor-1\geq t$ for $t\geq 3$ from Lemma 6. ∎
2 Notation
[…]
We write $e_{1},\ldots,e_{d}$ for the standard orthonormal basis vectors of $\mathbb{R}^{d}$. The orthogonal projection to the $i$-th coordinate in $\mathbb{R}^{d}$ is denoted by $\operatorname{proj}_{i}$ and by $\operatorname{proj}_{\mathbb{R}^{d-1}}$ we mean the orthogonal projection to the first $d-1$ coordinates.
Sets in $\mathbb{R}^{d}$.
Given $A\subset\mathbb{R}^{d}$ and $r>0$, we say that $A$ is $r$-separated if $\left\|a-a^{\prime}\right\|\geq r$ for every $a^{\prime}\in A,a\neq a^{\prime}$. We say that $A$ is an $r$-net if for every $x\in\mathbb{R}^{d}$ there is $a\in A$ such that $\left\|x-a\right\|\leq r$. Whenever there are $r,s>0$ so that $A$ is $s$-separated $r$-net, we call $A$ a separated net.
We write $B(x,r)$ and $\overline{B}(x,r)$ respectively for the open and closed euclidean balls with centre $x\in\mathbb{R}^{d}$ and radius $r\geq 0$. We also use the same notation for neighbourhoods of sets, i.e, $B(A,r):=\bigcup_{x\in A}B(x,r)$, where $A\subseteq\mathbb{R}^{d}$, and similarly for $\overline{B}(A,r)$.
Given $A\subseteq\mathbb{R}^{d}$, by $\partial A,\operatorname{Int}A$ and $\overline{A}$ we denote the boundary, the interior and the closure of $A$ in $\mathbb{R}^{d}$, respectively.
Mappings.
For a mapping $f\colon A^{\prime}\subseteq\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $A\subseteq A^{\prime}$, we write $f|_{A}$ for the restriction of $f$ to $A$. The notation $\operatorname{dom}(f),\operatorname{image}(f)$ refer to the domain and the image of $f$, respectively.
For $A\subseteq\mathbb{R}^{d}$ and a mapping $f\colon A\to\mathbb{R}^{n}$ we let
$\operatorname{Lip}(f):=\sup\left\{\frac{\left\|f(y)-f(x)\right\|}{\left\|y-x\right\|}\colon x,y\in\operatorname{dom}(f),x\neq y\right\}.$
In the case that $f$ is injective, we further define
$\operatorname{bilip}(f):=\max\left\{\operatorname{Lip}(f),\operatorname{Lip}(f^{-1})\right\}.$
Given $L\geq 1$, we say that $f$ is $L$-Lipschitz if $\operatorname{Lip}(f)\leq L$ and that it is $L$-bilipschitz if $\operatorname{bilip}(f)\leq L$. Further, we say that $f$ is Lipschitz (bilipschitz) if there is $L<\infty$ such that $f$ is $L$-Lipschitz ($L$-bilipschitz).
3 Bilipschitz Swapping and Gluing.
The main aim of the present section is to prove the following theorem, which provides a bilipschitz mapping that efficiently performs ‘simultaneous swappings’ of a family of specified pairs $(x,y_{x})\in\mathbb{R}^{d}\times\mathbb{R}^{d}$.
Theorem 3.1.
Let $X\subseteq\mathbb{R}^{d}$. For each $x\in X$ let $y_{x}\in\mathbb{R}^{d}$, $r_{x}$ be a positive real number and $\mathfrak{U}_{x}:=B([x,y_{x}],r_{x})$. Suppose that $(\mathfrak{U}_{x})_{x\in X}$ is pairwise disjoint,
$\sup_{x\in X,\,y_{x}\neq x}\frac{r_{x}}{\left\|y_{x}-x\right\|}\leq\frac{1}{2}\qquad\text{and}\qquad\sup_{x\in X,\,y_{x}\neq x}\frac{\left\|y_{x}-x\right\|^{2}}{r_{x}^{2}}<\infty.$
Then there is a bilipschitz mapping $\Gamma\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that
  1. (i)
    $\Gamma(x)=y_{x}$ and $\Gamma(y_{x})=x$ for all $x\in X$.
  2. (ii)
    $\Gamma(p)=p$ for all $p\in\mathbb{R}^{d}\setminus\bigcup_{x\in X}\mathfrak{U}_{x}$.
  3. (iii)
    $\operatorname{bilip}(\Gamma)\leq\max\left\{1,\sup_{x\in X,\,y_{x}\neq x}\frac{4\left\|y_{x}-x\right\|^{2}}{r_{x}^{2}}\right\}$.
The ‘building blocks’ of the mapping $\Gamma$ from Theorem 3.1 will be bilipschitz mappings which perform a single swap of a specified pair $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$. So our first step towards the proof of Theorem 3.1 will be the construction of such mappings $\tau$ in the next Lemmas 1 and 1. We will then need an important ‘gluing lemma’, Lemma 1, to glue a family of ‘single swap mappings’ $\tau$ together to form $\Gamma$. Lemma 1 is also useful beyond the proof of Theorem 3.1.
Lemma 1.
Let $d\in\mathbb{N}$, $d\geq 2$ and for $\theta\in\mathbb{R}$ let $R_{\theta}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ be the linear mapping with matrix
$R_{\theta}=\begin{pmatrix}\cos\theta&-\sin\theta&0&0&\ldots&0\\ \sin\theta&\cos\theta&0&\ddots&\ddots&0\\ 0&0&1&0&\ddots&0\\ 0&0&0&1&\ddots&0\\ 0&0&0&0&\ddots&0\\ 0&0&0&0&0&1\end{pmatrix}.$
Here we use the same notation for the linear mapping and its matrix.
  1. (i)
    Let $\theta_{1},\theta_{2}\in\mathbb{R}$ and $y\in\mathbb{R}^{d}$. Then
    $\left\|R_{\theta_{2}}(y)-R_{\theta_{1}}(y)\right\|\leq\left|\theta_{2}-\theta_{1}\right|\left\|y\right\|.$
  2. (ii)
    Let $t_{0}>0$ and $\psi\colon[0,\infty)\to\mathbb{R}$ be a Lipschitz mapping with $\psi(t)=0$ for all $t\geq t_{0}$. Let $\Phi=\Phi(\psi)\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ be defined by
    $\Phi(x)=R_{\psi(\left\|x\right\|)}(x).$
    Then $\Phi$ is $\left(\operatorname{Lip}(\psi)t_{0}+1\right)$-bilipschitz and $\Phi(x)=x$ for all $x\in\mathbb{R}^{d}\setminus B(0,t_{0})$.
Proof.
  1. (i)
    For $\alpha\in\mathbb{R}$, let $\rho_{\alpha}\colon\mathbb{R}^{2}\to\mathbb{R}^{2}$ denote the anticlockwise rotation around the origin in $\mathbb{R}^{2}$ through angle $\alpha$. Then, we observe that
    $R_{\theta}(x_{1},\ldots,x_{d})=(\rho_{\theta}(x_{1},x_{2}),x_{3},\ldots,x_{d})$
    for all $\theta\in\mathbb{R}$ and $x=(x_{1},\ldots,x_{d})\in\mathbb{R}^{d}$. Therefore,
    $\left\|R_{\theta_{2}}(y)-R_{\theta_{1}}(y)\right\|=\left\|\rho_{\theta_{2}}(y_{1},y_{2})-\rho_{\theta_{1}}(y_{1},y_{2})\right\|\\ \leq\left|\theta_{2}-\theta_{1}\right|\left\|(y_{1},y_{2})\right\|\leq\left|\theta_{2}-\theta_{1}\right|\left\|y\right\|.$
  2. (ii)
    First, note that $\Phi(\psi)\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a bijection with $\Phi(\psi)^{-1}=\Phi(-\psi)$ and $\Phi(\psi)(x)=x$ for all $x\in\mathbb{R}^{d}\setminus B(0,t_{0})$. Letting $\zeta\in\left\{-\psi,+\psi\right\}$ it remains to verify that $\Phi(\zeta)$ is $\left(\operatorname{Lip}(\psi)t_{0}+1\right)$-Lipschitz. Since $\Phi(\zeta)$ coincides with the identity outside of $B(0,t_{0})$ it suffices to verify the Lipschitz bound between pairs of points where at least one point lies in $B(0,t_{0})$. Let $x\in\mathbb{R}^{d}$ and $y\in B(0,t_{0})$. Then applying part (i) we get
    $\left\|\Phi(\zeta)(y)-\Phi(\zeta)(x)\right\|\leq\left\|R_{\zeta(\left\|y\right\|)}(y)-R_{\zeta(\left\|x\right\|)}(y)\right\|+\left\|R_{\zeta(\left\|x\right\|)}(y-x)\right\|\\ \leq\left|\zeta(\left\|y\right\|)-\zeta(\left\|x\right\|)\right|t_{0}+\left\|y-x\right\|\\ \leq\operatorname{Lip}(\zeta)\left|\left\|y\right\|-\left\|x\right\|\right|t_{0}+\left\|y-x\right\|\leq\left(\operatorname{Lip}(\zeta)t_{0}+1\right)\left\|y-x\right\|.\qed$
[…]
Proof.
The statement is invariant under scaling, so we may assume that $\left\|y-x\right\|=1$ and replace $r$ with $\eta:=\frac{r}{\left\|y-x\right\|}\leq\frac{1}{2}$. We may now also assume that $x=\frac{1}{2}e_{1}$ and $y=-\frac{1}{2}e_{1}$. We introduce a linear isomorphism $T\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ which maps the largest ellipsoid $E$ inside $B([x,y],\eta)$ centred at 0 and with axes parallel to the coordinate axes to a ball $T(E)=B(0,R)$ of some radius $R>1$ so that $Tx=e_{1}$ and $Ty=-e_{1}$. The last condition tells us that $Te_{1}=2e_{1}$, which then implies that $R=2\left(\frac{1}{2}+\eta\right)$. For $i=2,3,\ldots,d$, we then have that $T(\eta e_{i})=Re_{i}$. Thus, we get that $T$ is defined by
$Te_{i}=\begin{cases}2e_{1}&\text{if }i=1,\\ \left(2+\frac{1}{\eta}\right)e_{i}&\text{if }i=2,3,\ldots,d.\end{cases}$
Note that
$T^{-1}\left(B\left(0,1+2\eta\right)\right)\subseteq B([x,y],\eta),\qquad\operatorname{Lip}(T)=2+\frac{1}{\eta},\qquad\operatorname{Lip}(T^{-1})=\frac{1}{2}.$
We now define $\psi\colon[0,\infty)\to\mathbb{R}$ by
$\psi(t)=\begin{cases}\pi&\text{if }0\leq t\leq 1,\\ \frac{\left(1+2\eta-t\right)\pi}{2\eta}&\text{if }1\leq t\leq 1+2\eta,\\ 0&\text{if }t\geq 1+2\eta.\end{cases}$
and note that $\operatorname{Lip}(\psi)=\frac{\pi}{2\eta}$. The desired mapping may now be defined as $\tau:=T^{-1}\circ\Phi(\psi)\circ T$, where $\Phi(\psi)$ is given by the conclusion of Lemma 1 applied with $t_{0}=1+2\eta$, so that
$\operatorname{bilip}\left(\Phi(\psi)\right)\leq\frac{\pi}{2\eta}\cdot\left(1+2\eta\right)+1\leq\frac{\pi}{\eta}+1,$
Multiplying $\operatorname{bilip}\left(\Phi(\psi)\right)$ by $\operatorname{Lip}(T^{-1})\operatorname{Lip}(T)\leq\frac{1}{\eta}$ we obtain $\operatorname{bilip}(\tau)\leq\frac{\pi}{\eta^{2}}+\frac{1}{\eta}\leq\frac{4}{\eta^{2}}$. With the properties of $\Phi(\psi)$ given by Lemma 1, property (i) of $\tau$ follows from
$T\left(\mathbb{R}^{d}\setminus B([x,y],r)\right)\subseteq\mathbb{R}^{d}\setminus B(0,1+2\eta).$
Moreover, the calculation
$\tau(x)=T^{-1}\circ\Phi(\psi)\circ T(x)=T^{-1}\left(R_{\psi(1)}(e_{1})\right)=T^{-1}\left(R_{\pi}(e_{1})\right)=T^{-1}(-e_{1})=y,$
and a similar one for $\tau(y)$ verify property (ii) of $\tau$. ∎
[…]
Proof.
We first show that $f$ is injective. Let $x\in\overline{A_{i}}$ and $y\in\overline{A_{j}}$ for some $i,j\in I$ such that $f(x)=f(y)$. If $i=j$, we use that $f|_{\overline{A_{i}}}$ is bilipschitz to deduce $x=y$. So assume $i\neq j$. Then, the assumptions $A_{i}\cap A_{j}=\emptyset$ and $f(\overline{A_{k}})=\overline{A_{k}}$ for $k=i,j$ imply $f(x)=f(y)\in\partial A_{i}\cap\partial A_{j}$. By Brouwer’s Invariance of Domain [11, Thm 2B.3], $\partial f(A_{k})=f(\partial A_{k})$ for every $k\in I$. Thus, $x\in\partial A_{i}$ and $y\in\partial A_{j}$. However, $f|_{\bigcup_{i\in I}\partial A_{i}}$ is bilipschitz as well, and thus, $x=y$. Consequently, $f^{-1}$ is well-defined.
Since $f(A_{i})=A_{i}$ for every $i\in I$ and as $f(\bigcup_{i\in I}\partial A_{i})\supseteq\bigcup_{i\in I}f(\partial A_{i})=\bigcup_{i\in I}\partial f(A_{i})$, we see that the assumptions of the lemma are satisfied for $f^{-1}$ and $\left(f(A_{i})\right)_{i\in I}$ instead of $f$ and $(A_{i})_{i\in I}$. Therefore, it suffices to show that $f$ is $C$-Lipschitz.
Let $i,j\in I,i\neq j$, and $x\in A_{i}\setminus\overline{A_{j}}$ and $y\in A_{j}\setminus\overline{A_{i}}$. We set $z_{1}:=x$, $z_{4}:=y$ and take $z_{2}$ as any point of $[x,y]\cap\partial A_{i}$. Since $x\in A_{i}$ and $y\notin\overline{A_{i}}$, $z_{2}$ is well-defined. Note that $\partial A_{i}\cap A_{j}=\emptyset$, so $z_{2}\notin A_{j}$. Finally, we take $z_{3}$ as any point of $[z_{2},y]\cap\partial A_{j}$. Therefore, by the assumptions on the $C$-bilipschitz property,
$\left\|f(y)-f(x)\right\|\leq\sum_{k=2}^{4}\left\|f(z_{k})-f(z_{k-1})\right\|\leq C\sum_{k=2}^{4}\left\|z_{k}-z_{k-1}\right\|=C\left\|y-x\right\|.\qed$
We now have all the necessary tools to prove Theorem 3.1.
Proof of Theorem 3.1.
For each $x\in X$ with $y_{x}=x$ let $\tau_{x}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ be the identity mapping. For each remaining $x\in X$ let $\tau_{x}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ be the mapping given by Lemma 1 applied to $x$, $y_{x}$ and $r_{x}$. Setting $C:=\displaystyle\max\left\{1,\sup_{x\in X,\,y_{x}\neq x}\frac{4\left\|y_{x}-x\right\|^{2}}{r_{x}^{2}}\right\}$, note that each $\tau_{x}$ is $C$-bilipschitz and equal to the identity outside $\mathfrak{U}_{x}$. We then define $\Gamma\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ by
$\Gamma(p)=\begin{cases}\tau_{x}(p)&\text{if $x\in X$ and $p\in\mathfrak{U}_{x}$},\\ p&\text{if }p\in\mathbb{R}^{d}\setminus\bigcup_{x\in X}\mathfrak{U}_{x}.\end{cases}$
Obviously, $\Gamma$ and $\Gamma^{-1}$ are well-defined. Due to the symmetry between $\Gamma$ and $\Gamma^{-1}$, it is enough to show that $\Gamma$ is $C$-Lipschitz.
Let $x_{1},x_{2}\in X$, $x_{1}\neq x_{2}$. Given any $y_{i}\in\mathfrak{U}_{x_{i}}$, $i\in[2]$, we take $z_{i}\in[y_{1},y_{2}]\cap\partial\mathfrak{U}_{x_{i}}$. The disjointness of the collection $\left(\mathfrak{U}_{x}\right)_{x\in X}$ implies that $z_{1},z_{2}\notin\bigcup_{x\in X}\mathfrak{U}_{x}$. By the triangle inequality
$\left\|\Gamma(y_{1})-\Gamma(y_{2})\right\|$ $\leq\left\|\Gamma(z_{1})-\Gamma(z_{2})\right\|+\sum_{i=1}^{2}\left\|\Gamma(y_{i})-\Gamma(z_{i})\right\|$
$\leq\left\|z_{1}-z_{2}\right\|+C\sum_{i=1}^{2}\left\|y_{i}-z_{i}\right\|\leq C\left\|y_{1}-y_{2}\right\|.$
Similarly, when $y_{1}\in\mathfrak{U}_{x_{1}}$ and $y_{2}\notin\bigcup_{x\in X}\mathfrak{U}_{x}$, we can set $z_{2}=y_{2}$ in the above chain of inequalities. We conclude that $\Gamma$ is $C$-Lipschitz. ∎
4 Reduction to the integer lattice.
This section is dedicated to the full proof of Theorem 1.1, in which we reduce the bilipschitz extension problem for all separated nets to that for the integer lattice. Before proceeding, the reader may wish to read the sketch proof of Theorem 1.1 in the introduction.
Lemma 2.
Let $d\in\mathbb{N}$, $d\geq 2$, $r>0$ and $A\subseteq\mathbb{R}^{d}$ be $r$-separated. Then there exists a $16\max\left\{\frac{3d}{r},1\right\}$-bilipschitz mapping $\Phi\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $\Phi(A)\subseteq\mathbb{Z}^{d}$.
Proof.
[…]
$r_{x}:=\begin{cases}\left\|y_{x}-\Psi(x)\right\|/2&\text{if }y_{x}\neq\Psi(x),\\ \sqrt{d}&\text{otherwise.}\end{cases}$
and observe that $\mathfrak{U}_{x}:=B([\Psi(x),y_{x}],r_{x})\subseteq B(y,\sqrt{d})$. Hence $(\mathfrak{U}_{x})_{x\in A}$ is pairwise disjoint. Let $\Pi\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ be the 16-bilipschitz mapping given by the application of Theorem 3.1 to $X=\Psi(A)$, $\Psi(x)$ (in place of $x$), $y_{x}$ and $r_{x}$ for each $x\in A$. All that remains is to set $\Phi=\Pi\circ\Psi$. ∎
Lemma 3.
Let $d\in\mathbb{N}$, $d\geq 2$, $L\geq 1$, $X\subseteq\mathbb{Z}^{d}$ and $f\colon X\to\mathbb{R}^{d}$ be an $L$-bilipschitz mapping. Suppose that there exists $\lambda\geq 1$ such that $\mathbb{R}^{d}\subseteq\bigcup_{x\in X}\overline{B}(x,\lambda)$. Then there exists a bilipschitz extension $F\colon\mathbb{Z}^{d}\to\mathbb{R}^{d}$ of $f$ with
$\operatorname{Lip}(F)\leq 4\lambda L,\qquad\operatorname{Lip}(F^{-1})\leq 24\lambda^{2}L\sqrt{d}.$
Proof.
For each $y\in\mathbb{Z}^{d}\setminus X$ choose $\alpha_{y}\in X$ such that $y\in\overline{B}(\alpha_{y},\lambda)$. Set $\alpha_{x}=x$ for every $x\in X$. For any $t\geq\sqrt{d}$, comparing a euclidean ball $\overline{B}(0,t)$ with cubes $[-t,t]^{d}$ and $\left[-\frac{t}{\sqrt{d}},\frac{t}{\sqrt{d}}\right]^{d}$ we get the following bounds:
$2^{d}\left(\frac{t}{\sqrt{d}}-1\right)^{d}\leq\left|\overline{B}(0,t)\cap\mathbb{Z}^{d}\right|\leq 2^{d}(t+1)^{d}$ (1)
In particular, for each $x\in X$ we have
|$\left|\left\{y\in\mathbb{Z}^{d}\colon\alpha_{y}=x\right\}\right|\leq\left|\overline{B}(0,\lambda)\cap\mathbb{Z}^{d}\right|\leq 2^{d}(\lambda+1)^{d}.$
Next, observe that the balls $(B(f(x),\frac{1}{2L}))_{x\in X}$ are pairwise disjoint. Setting $N:=12\lambda\sqrt{d}L$ and using (1) we infer that for every $w\in\mathbb{R}^{d}$
$\left|\left(\overline{B}\left(w,\frac{1}{4L}\right)\right)\cap\left(w+\frac{1}{N}\mathbb{Z}^{d}\right)\right|=\left|\overline{B}\left(0,\frac{N}{4L}\right)\cap\mathbb{Z}^{d}\right|\geq 2^{d}(\lambda+1)^{d}.$
As $x\in\left\{y\in\mathbb{Z}^{d}\colon\alpha_{y}=x\right\}$ for all $x\in X$, it follows that there is an injective mapping $F\colon\mathbb{Z}^{d}\to\mathbb{R}^{d}$ such that
$F(y)\in\left(\overline{B}\left(f(\alpha_{y}),\frac{1}{4L}\right)\right)\cap\left(f(\alpha_{y})+\frac{1}{N}\mathbb{Z}^{d}\right)\qquad\text{ for every $y\in\mathbb{Z}^{d}$}$
and $F(x)=f(x)$ for all $x\in X$. These conditions imply $\left\|F(p)-F(\alpha_{p})\right\|\leq 1/(4L)$ for every $p\in\mathbb{Z}^{d}$.
Let $p,q\in\mathbb{Z}^{d}$ with $p\neq q$. We verify the bilipschitz bounds on $\left\|F(q)-F(p)\right\|$. The definition of $(\alpha_{y})_{y\in\mathbb{Z}^{d}}$ implies the following inequality, which we will use several times:
$\left|\left\|\alpha_{q}-\alpha_{p}\right\|-\left\|q-p\right\|\right|\leq 2\lambda.$
Since $\left\|q-p\right\|\geq 1$ we have
$\left\|F(q)-F(p)\right\|\leq\left\|F(\alpha_{q})-F(\alpha_{p})\right\|+\frac{1}{2L}\leq\left(L+2L\lambda+\frac{1}{2L}\right)\left\|q-p\right\|.$
The lower bilipschitz bound is easy in the case that $\alpha_{p}\neq\alpha_{q}$: then it holds that $\left\|F(q)-F(p)\right\|\geq\left\|F(\alpha_{q})-F(\alpha_{p})\right\|-\frac{1}{2L}\geq\frac{1}{2L}$. So assume now that $\alpha_{p}=\alpha_{q}$. Then $\left\|q-p\right\|\leq 2\lambda$ and $\left\|F(q)-F(p)\right\|\geq 1/N$, so $\left\|F(q)-F(p)\right\|\geq\frac{1}{2\lambda N}\left\|q-p\right\|$. This completes the verification of the bilipschitz property for $F$ and substituting in the set value $N=12\lambda\sqrt{d}L$ to the estimates above delivers the desired bounds on $\operatorname{Lip}(F)$ and $\operatorname{Lip}(F^{-1})$. ∎
Proof of Theorem 1.1.
For the equivalence, it suffices to prove the implication (i)$\Rightarrow$(ii), as the other one is trivial. Assume that (i) holds, let $A\subset\mathbb{R}^{d}$ be an $r$-separated $R$-net and let $f\colon A\to\mathbb{R}^{d}$ be bilipschitz. We set $K:=16\max\left\{\frac{3d}{r},1\right\}$ and apply Lemma 2 to $A$ to get a $K$-bilipschitz mapping $\Phi\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $\Phi(A)\subset\mathbb{Z}^{d}$. Next, we apply Lemma 3 to $f\circ\Phi^{-1}|_{\Phi(A)}$ and obtain its bilipschitz extension $g\colon\mathbb{Z}^{d}\to\mathbb{R}^{d}$ with $\operatorname{bilip}(g)\leq 24\sqrt{d}R^{2}K^{3}\operatorname{bilip}(f)$. Finally, (i) provides us with a bilipschitz extension $G\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ of $g$. The desired extension of $f$ is then given by $F:=G\circ\Phi$. If (i) holds with $\operatorname{bilip}(F)\leq C_{d}(\operatorname{bilip}(f))$ for some monotone function $C_{d}\colon[1,\infty)\to[1,\infty)$ then the bilipschitz extension $F$ of $f\colon A\to\mathbb{R}^{d}$ constructed above to verify (ii) has bilipschitz constant bounded above by $\operatorname{bilip}(\Phi)\operatorname{bilip}(G)\leq K\cdot C_{d}\left(24\sqrt{d}R^{2}K^{3}\operatorname{bilip}(f)\right)$. ∎
5 Bilipschitz permutations of grid points on a hyperplane.
The main result of the present section provides a bilipschitz transformation of a slab $\mathbb{R}^{d-1}\times[m-1/2,m+1/2]$, for some $m\in\mathbb{R}$, that extends a given permutation of $\frac{1}{N}\mathbb{Z}^{d-1}\times\left\{m\right\}$ with a finite distance to the identity; this is one of the main steps in the proof of Theorem 1.2.
Theorem 5.1.
Let $d\geq 2$. Then there exists a monotone increasing function $\beta_{d}\colon[1,\infty)\to[1,\infty)$ such that $2\lfloor t\rfloor-1\leq\beta_{d}(t)\leq\exp(8dt)$ for all $t\in[1,\infty)$ and the following statement holds:
Let $T,N\in\mathbb{N}$, $m\in\mathbb{R}$ and $\sigma\colon\frac{1}{N}\mathbb{Z}^{d-1}\to\frac{1}{N}\mathbb{Z}^{d-1}$ be a permutation with $\left\|\sigma(x)-x\right\|\leq T$ for all $x\in\frac{1}{N}\mathbb{Z}^{d-1}$. Then there exists a
$N\beta_{d}(6NT)^{3^{d-1}}$bilipschitz mapping
$\Upsilon\colon\mathbb{R}^{d-1}\times\left[m-\frac{1}{2},m+\frac{1}{2}\right]$ $\to\mathbb{R}^{d-1}\times\left[m-\frac{1}{2},m+\frac{1}{2}\right]$
such that
  1. (i)
    $\Upsilon(x)=x$ for all $x\in\mathbb{R}^{d-1}\times\left\{m-\frac{1}{2},m+\frac{1}{2}\right\}$.
  2. (ii)
    $\Upsilon(x,m)=(\sigma(x),m)$ for all $x\in\frac{1}{N}\mathbb{Z}^{d-1}$.
Moreover, the function $\beta_{d}$ may be defined as in Lemma 6.
Before we can prove Theorem 5.1, we need to make a series of intermediate steps.
[…]
Lemma 4.
Let $l,T\in\mathbb{N}$, the set $\widehat{Q}\subseteq\mathbb{R}^{l}$ be defined by
$\widehat{Q}:=\left(-\frac{1}{4},\ldots,-\frac{1}{4}\right)+[-T,2T]^{l},$
and $\phi\colon\mathbb{Z}^{l}\to\mathbb{Z}^{l}$ be a permutation such that $\left\|\phi(x)-x\right\|\leq T$ for all $x\in\mathbb{Z}^{l}$. Then there exist permutations $\sigma_{1},\sigma_{2},\ldots,\sigma_{3^{l}}\colon\frac{1}{2}\mathbb{Z}^{l}\to\frac{1}{2}\mathbb{Z}^{l}$ such that
  1. (i)
    For each $k\in[3^{l}]$ there exists $p_{k}\in\left\{-1,0,1\right\}^{l}$ such that for every $z\in\mathbb{Z}^{l}$ the restriction $\sigma_{k}|_{\frac{1}{2}\mathbb{Z}^{l}\cap\left(\widehat{Q}+T(3z+p_{k})\right)}$ is a permutation of $\frac{1}{2}\mathbb{Z}^{l}\cap\left(\widehat{Q}+T\left(3z+p_{k}\right)\right)$.
  2. (ii)
    $\phi=\left(\sigma_{3^{l}}\circ\sigma_{3^{l}-1}\circ\ldots\sigma_{1}\right)|_{\mathbb{Z}^{l}}$.
Proof.
Let $Q:=(-1/4,\ldots,-1/4)+[0,T]^{l}$ and $\Delta\colon\mathbb{Z}^{l}\to\mathbb{Z}^{l}$ be the mapping uniquely defined by the condition
$\phi(x)\in Q+T\Delta(x)\qquad\text{for all }x\in\mathbb{Z}^{l}.$ (2)
Moreover, let $\left\{p_{1},\ldots,p_{3^{l}}\right\}$ be an enumeration of $\left\{-1,0,1\right\}^{l}$. As $\phi$ is a permutation of $\mathbb{Z}^{l}$, for every $z\in\mathbb{Z}^{l}$,
$\left|\frac{1}{2}\mathbb{Z}^{l}\cap(Q+Tz)\right|-\left|\mathbb{Z}^{l}\cap(Q+Tz)\right|\geq\left|\Delta^{-1}(\left\{z\right\})\right|.$ (3)
The idea is to first define preliminary version of permutations $\sigma_{i},i\in[3^{l}]$, satisfying (i), but instead of (ii) they will only ensure that $x\in\mathbb{Z}^{l}$ is mapped inside $Q+T\Delta(x)$ using the ‘available spots’ in $\frac{1}{2}\mathbb{Z}^{l}\setminus\mathbb{Z}^{l}$. Then it suffices to permute the points of $\frac{1}{2}\mathbb{Z}^{l}$ inside each $Q+Tz,z\in\mathbb{Z}^{l},$ so that each $x$ is mapped to $\phi(x)$.
By (3), for every $i\in[3^{l}]$ and $w\in 3\mathbb{Z}^{l}$ there exists an injective mapping
$\iota_{w}^{i}\colon\left\{x\in\Delta^{-1}(\left\{w+p_{i}\right\})\colon x\notin Q+T(w+p_{i})\right\}\to\left(\frac{1}{2}\mathbb{Z}^{l}\setminus\mathbb{Z}^{l}\right)\cap\left(Q+T(w+p_{i})\right).$ (4)
For every $z\in\mathbb{Z}^{l}$, whenever $x\in\Delta^{-1}(\left\{z\right\})$, it holds that $\phi(x)\in Q+Tz$ by (2), which in turn implies $x\in\widehat{Q}+Tz$ by the assumption on $\phi$. Thus,
$\operatorname{dom}(\iota_{w}^{i})\subseteq\mathbb{Z}^{l}\cap(\widehat{Q}+T(w+p_{i})).$ (5)
It follows immediately from the definition that for every $i,j\in[3^{l}]$ and every $w,w^{\prime}\in 3\mathbb{Z}^{l}$ such that $(i,w)\neq(j,w^{\prime})$
$\operatorname{dom}(\iota_{w}^{i})\cap\operatorname{dom}(\iota_{w^{\prime}}^{j})=\emptyset,\quad\operatorname{image}(\iota_{w}^{i})\cap\operatorname{image}(\iota_{w^{\prime}}^{j})=\emptyset\quad\text{and}\quad\operatorname{dom}(\iota_{w}^{i})\cap\operatorname{image}(\iota_{w^{\prime}}^{j})=\emptyset.$ (6)
Now we can define bijections $\tilde{\sigma}_{i}\colon\frac{1}{2}\mathbb{Z}^{l}\to\frac{1}{2}\mathbb{Z}^{l}$ for every $i\in[3^{l}]$. For each $w\in 3\mathbb{Z}^{l}$ we set $\tilde{\sigma}_{i}|_{\frac{1}{2}\mathbb{Z}^{l}\cap(\widehat{Q}+T(w+p_{i}))}$ to be the permutation of $\frac{1}{2}\mathbb{Z}^{l}\cap(\widehat{Q}+T(w+p_{i}))$ which interchanges $x$ and $\iota_{w}^{i}(x)$ for each $x\in\operatorname{dom}(\iota_{w}^{i})$ and fixes all other points of $\frac{1}{2}\mathbb{Z}^{l}\cap(\widehat{Q}+T(w+p_{i}))$.
The crucial properties following from this definition are:
  1. (A)
    By (5), each $\tilde{\sigma}_{i}$ is a permutation of $\frac{1}{2}\mathbb{Z}^{l}$ that only shuffles points inside each tile of the tiling $\widehat{Q}+T(3\mathbb{Z}^{l}+p_{i})$.
  2. (B)
    By (6), for every $i,j\in[3^{l}],i\neq j$, and every $x\in\frac{1}{2}\mathbb{Z}^{l}$, if $\tilde{\sigma}_{i}(x)\neq x$, then $\tilde{\sigma}_{j}(x)=x$.
  3. (C)
    By (4), for every $x\in\mathbb{Z}^{l}$ such that $x\in Q+T\Delta(x)$ and every $i\in[3^{l}]$ it holds that $\tilde{\sigma}_{i}(x)=x$.
  4. (D)
    Since for every $z\in\mathbb{Z}^{l}$ there are unique $w\in 3\mathbb{Z}^{l}$ and $i\in[3^{l}]$ such that $z=w+p_{i}$, we infer that for every $x\in\mathbb{Z}^{l}$ such that $x\notin Q+T\Delta(x)$ there is unique $i\in[3^{l}]$ such that $\tilde{\sigma}_{i}(x)\neq x$; moreover, this $i$ satisfies $\tilde{\sigma}_{i}(x)\in Q+T\Delta(x)$.
Consequently, writing $\tilde{\sigma}:=\tilde{\sigma}_{3^{l}}\circ\cdots\circ\tilde{\sigma_{1}}$, (B)–(D) imply that $\tilde{\sigma}(x)\in Q+T\Delta(x)$ for every $x\in\mathbb{Z}^{l}$. Therefore, there exists a permutation $\tau\colon\frac{1}{2}\mathbb{Z}^{l}\to\frac{1}{2}\mathbb{Z}^{l}$ such that for each $z\in\mathbb{Z}^{l}$ the restriction $\tau|_{\frac{1}{2}\mathbb{Z}^{l}\cap\left(Q+Tz\right)}$ is a permutation of $\frac{1}{2}\mathbb{Z}^{l}\cap\left(Q+Tz\right)$ and $\tau\circ\tilde{\sigma}(x)=\phi(x)$ for each $x\in\mathbb{Z}^{l}$. Note that, trivially, for each $w\in 3\mathbb{Z}^{l}$ the restriction of $\tau$ to $\frac{1}{2}\mathbb{Z}^{l}\cap\left(\widehat{Q}+T(w+p_{3^{l}})\right)$ is a permutation of $\frac{1}{2}\mathbb{Z}^{l}\cap\left(\widehat{Q}+T(w+p_{3^{l}})\right)$. The theorem is now verified by the permutations $\sigma_{1},\sigma_{2},\ldots,\sigma_{3^{l}}\colon\frac{1}{2}\mathbb{Z}^{l}\to\frac{1}{2}\mathbb{Z}^{l}$ defined by
$\sigma_{i}:=\begin{cases}\tilde{\sigma}_{i}&\text{if }1\leq i\leq 3^{l}-1,\\ \tau\circ\tilde{\sigma}_{3^{l}}&\text{if }i=3^{l}.\end{cases}\qed$
Our bilipschitz realisation of permutations of cubical portions of $\mathbb{Z}^{l}$ is based on a decomposition of the permutation into a sequence of grid-adjacent transpositions, which are then realised using Lemma 1. In order to keep the bilipschitz constant of the final mapping as low as possible, we take advantage of the fact that transpositions of disjoint pairs of grid-adjacent points can be performed simultaneously without further increase in the bilipschitz constant. Thus, we want to decompose any given permutation into a small number of sets of pairwise disjoint grid-adjacent transpositions.
Lemma 5 (Attributed to Slepian in [6]; also see [4, Thm. 1]).
Let $l,S\in\mathbb{N}$. Every permutation $\sigma\colon[S]^{l}\to[S]^{l}$ can be decomposed as $\sigma=\sigma_{(2l-1)S}\circ\cdots\circ\sigma_{1}$, where $\sigma_{i}\colon[S]^{l}\to[S]^{l}$ is a permutation such that $\sigma_{i}=\sigma_{i}^{-1}$ and $\left\|x-\sigma_{i}(x)\right\|\leq 1$ for every $x\in[S]^{l}$ and $i\in[(2l-1)S]$.
This result has been attributed to Slepian in [6] in the context of routing and sorting on networks. The graph-theoretic perspective on the statement is perhaps the easiest to understand; we refer the reader to [9] for the graph-theoretic notions appearing in the following discussion.
Abstractly, the problem has been studied under the term of ‘routing on graphs’; given a graph $G=(V,E)$, a set of pebbles $\left\{p_{v}\colon v\in V\right\}$, one at each vertex of $G$, and a bijection $\tau\colon V\to V$, the goal is to find a sequence of matchings $P_{1},\ldots,P_{k}\subset E$ with the following property: At the step $i\in[k]$ for every $(u,v)\in P_{i}$ we switch the pebbles currently placed at $u$ and $v$. After the last step, every pebble $p_{v}$ is at the vertex $\tau(v)$ for $v\in V$. The goal is to minimize $k=k(\tau)$.
The routing number $\operatorname{rt}(G)$ of a graph $G$ is the maximum of $k(\tau)$ over all permutations $\tau\colon V\to V$. The graph relevant for Lemma 5 is the $l$-fold Cartesian product $P_{S}\,\square\cdots\square\,P_{S}$, where $P_{S}$ denotes the path on $S$ vertices; its vertices are naturally identified with $[S]^{l}$ and its edges then correspond to grid-adjacent pairs of points of $[S]^{l}$.
It is classical that $\operatorname{rt}(P_{S})=S$ (e.g., [13, 5.3.4.Ex.37]). The key observation is that $\operatorname{rt}(G\,\square\,H)\leq 2\operatorname{rt}(G)+\operatorname{rt}(H)$ holds for any two graphs $G,H$; a short and clear presentation of the relevant arguments can be found in [3, Thm. 4].
In the next lemma, we deal with permutations of $\mathbb{Z}^{d-1}$ that respect a fixed regular tiling of $\mathbb{R}^{d-1}$, that is, for each tile, the images of the lattice points remain inside the same tile. The lemma establishes the existence of a bilipschitz mapping $\Upsilon\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ that realizes the given permutation of $\mathbb{Z}^{d-1}$ seen as lying inside the hyperplane $\mathbb{R}^{d-1}\times\left\{0\right\}$.
Our upper bound on the bilipschitz constant of $\Upsilon$ is exponential in the size of the tiles and $d$. When $d$ is treated as a constant, it is the only cause of the exponential increase in the bilipschitz constant of the extension in Theorem 1.2. We believe that much better bound may be possible. Therefore, instead of stating our current upper bound directly in Lemma 6, we have abstracted the relevant quantity into a function $\beta_{d}$, so that any future improvement on the upper bound on $\beta_{d}$ yields directly an improvement in Theorem 1.2 without any need to repeat large parts of proofs from the present paper.
Lemma 6.
Let $d\in\mathbb{N}$ with $d\geq 2$ and $\mathcal{F}_{d}$ denote the collection of all monotone increasing functions $\psi\colon[1,\infty)\to[1,\infty)$ for which the following statement ((*)6) holds:
  1. (*)
    Given any $S\in\mathbb{N}$, $Q:=\left(-\frac{1}{2},-\frac{1}{2},\ldots,-\frac{1}{2}\right)+\left[0,S\right]^{d-1}$ and any permutation $\sigma\colon\mathbb{Z}^{d-1}\to\mathbb{Z}^{d-1}$ such that for each $z\in\mathbb{Z}^{d-1}$ the restriction $\sigma|_{\mathbb{Z}^{d-1}\cap\left(Q+Sz\right)}$ is a permutation of $\mathbb{Z}^{d-1}\cap\left(Q+Sz\right)$, there exists an $\psi(S)$-bilipschitz mapping $\Upsilon\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ with the following properties:
    1. (a)
      $\Upsilon(x)=x$ for all $x\in\mathbb{R}^{d-1}\times\left(\mathbb{R}\setminus\left(-\frac{1}{2},\frac{1}{2}\right)\right)$.
    2. (b)
      $\Upsilon(x,0)=(\sigma(x),0)$ for all $x\in\mathbb{Z}^{d-1}$.
Let $\beta_{d}\colon[1,\infty)\to[1,\infty)$ be the pointwise infimum of the collection $\mathcal{F}_{d}$. Then $\beta_{d}\in\mathcal{F}_{d}$ and $2\lfloor t\rfloor-1\leq\beta_{d}(t)\leq\exp\left(8dt\right)$ for all $t\in[1,\infty)$.
Proof.
We first show that $\exp(8dt)\in\mathcal{F}_{d}$. Fixing $S\in\mathbb{N}$, $Q$ and $\sigma\colon\mathbb{Z}^{d-1}\to\mathbb{Z}^{d-1}$ according to the hypothesis of ((*)6) we do this by constructing an $\exp(8dS)$-bilipschitz mapping $\Upsilon\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ with properties ((*)6) and ((*)6). Fix $z\in\mathbb{Z}^{d-1}$ and set $K:=(2d-3)S\leq 2dS$. Then $\Upsilon|_{\left(Q+Sz\right)\times\mathbb{R}}$ is defined as follows:
An application of Lemma 5 to $\sigma|_{\mathbb{Z}^{d-1}\cap\left(Q+Sz\right)}$ provides us with $K$ permutations $\sigma_{1},\ldots,\sigma_{K}$ of $\mathbb{Z}^{d-1}\cap\left(Q+Sz\right)$ decomposing $\sigma|_{\mathbb{Z}^{d-1}\cap\left(Q+Sz\right)}$. For each $k\in[K]$ and $x\in\mathbb{Z}^{d-1}\cap\left(Q+Sz\right)$, we have $\sigma_{k}(x)\in\mathbb{Z}^{d-1}\cap\left(Q+Sz\right)$ with $\left\|\sigma_{k}(x)-x\right\|\leq 1$. Moreover the sets in $\left\{\left\{x,\sigma_{k}(x)\right\}\colon x\in\mathbb{Z}^{d-1}\cap\left(Q+Sz\right)\right\}$ are pairwise disjoint for each $k\in[K]$. For each $k\in[K]$ let the 161616-bilipschitz mapping $\Gamma_{k}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ be given by Theorem 3.1 applied to $X=\left(\mathbb{Z}^{d-1}\cap\left(Q+Sz\right)\right)\times\left\{0\right\}\subseteq\mathbb{R}^{d}$, $y_{(x,0)}=(\sigma_{k}(x),0)\in\mathbb{R}^{d}$ and $r_{(x,0)}=1/2$ for each $(x,0)\in X$. Note that
$\Gamma_{k}((x,0))=(\sigma_{k}(x),0)$ for every $x\in\mathbb{Z}^{d-1}\cap\left(Q+Sz\right)$
and
$\Gamma_{k}(x)=x\qquad\text{for all }x\in\mathbb{R}^{d}\setminus\left(\operatorname{Int}Q+Sz\right)\times\left(-\frac{1}{2},\frac{1}{2}\right).$
The desired mapping $\Upsilon|_{\left(Q+Sz\right)\times\mathbb{R}}$ may now be defined by
$\Upsilon|_{\left(Q+Sz\right)\times\mathbb{R}}=\left(\Gamma_{K}\circ\Gamma_{K-1}\circ\cdots\circ\Gamma_{1}\right)|_{\left(Q+Sz\right)\times\mathbb{R}},$
so that
$\operatorname{bilip}(\Upsilon|_{\left(Q+Sz\right)\times\mathbb{R}})\leq\prod_{k=1}^{K}\operatorname{bilip}(\Gamma_{k})=16^{K}\leq\left(2^{4}\right)^{2dS}=\exp\left(8dS\right).$
Finally we apply Lemma 1 to $\Upsilon$ and the collection of open sets $((\operatorname{Int}Q+Sz)\times\mathbb{R})_{z\in\mathbb{Z}^{d-1}}$ to deduce $\operatorname{bilip}(\Upsilon)\leq\exp\left(8dS\right)$.
Having shown that $\exp\left(8dt\right)\in\mathcal{F}_{d}$ it follows immediately that $\beta_{d}\colon[1,\infty)\to[1,\infty)$ is well-defined and $\beta_{d}(t)\leq\exp\left(8dt\right)$ for all $t\in\mathbb{R}$. Moreover, $\beta_{d}$ is monotone increasing because any pointwise infimum of a non-empty collection of monotone increasing functions is such. It only remains to justify that statement ((*)6) holds with $\psi=\beta_{d}$. Fixing $S\in\mathbb{N}$ and a permutation $\sigma\colon\mathbb{Z}^{d-1}\to\mathbb{Z}^{d-1}$ of the type in ((*)6) we may choose for each $n\in\mathbb{N}$ a $\left(\beta_{d}(S)+\frac{1}{n}\right)$-bilipschitz $\Upsilon_{n}$ with properties ((*)6) and ((*)6). A standard application of the Arzelà–Ascoli Theorem provides a subsequence of $(\Upsilon_{n})_{n\in\mathbb{N}}$ converging uniformly on compact subsets of $\mathbb{R}^{d}$ to a $\beta_{d}(S)$-bilipschitz $\Upsilon$ inheriting ((*)6) and ((*)6).
For each $S\in\mathbb{N}$, taking $\sigma$ in ((*)6) to be the permutation of $\mathbb{Z}^{d-1}$ which swaps (0,,0)(0$\ldots$,0)( 0 , … , 0 ) and $(S-1,0,\ldots,0)$, $(S,0,\ldots,0)$ and $(2S-1,0,\ldots,0)$ and leaves all other points fixed shows that $\beta_{d}(S)\geq 2S-1$. The inequality $\beta_{d}(t)\geq 2\lfloor t\rfloor-1$ for all $t\in[1,\infty)$ now follows by the monotonicity of $\beta_{d}$. ∎
Proof of Theorem 5.1.
Let the function $\beta_{d}\colon[1,\infty)\to[1,\infty)$ be given by Lemma 6. Suppose that the theorem is valid whenever $N=1$ and $m=0$. Then we may consider the permutation $\widetilde{\sigma}\colon\mathbb{Z}^{d-1}\to\mathbb{Z}^{d-1}$ defined by $\widetilde{\sigma}(x)=N\sigma\left(\frac{x}{N}\right)$ and note that $\left\|\widetilde{\sigma}(x)-x\right\|\leq NT$ for all $x\in\mathbb{Z}^{d-1}$. Let $\widetilde{\Upsilon}\colon\mathbb{R}^{d-1}\times\left[-\frac{1}{2},\frac{1}{2}\right]\to\mathbb{R}^{d-1}\times\left[-\frac{1}{2},\frac{1}{2}\right]$ be the $\beta_{d}(6NT)^{3^{d-1}}$-bilipschitz mapping given by the version of the theorem with $N=1$ and $m=0$. The full statement is now verified by the mapping $\Upsilon\colon\mathbb{R}^{d-1}\times\left[m-\frac{1}{2},m+\frac{1}{2}\right]\to\mathbb{R}^{d-1}\times\left[m-\frac{1}{2},m+\frac{1}{2}\right]$ defined by
$\Upsilon(x)=\rho^{-1}\circ\widetilde{\Upsilon}\circ\rho(x)\qquad\text{for all $x\in\mathbb{R}^{d-1}\times\left[m-\frac{1}{2},m+\frac{1}{2}\right]$,}$
where $\rho\colon\mathbb{R}^{d-1}\times\left[m-\frac{1}{2},m+\frac{1}{2}\right]\to\mathbb{R}^{d-1}\times\left[-\frac{1}{2},\frac{1}{2}\right]$ is the affine bijection given by $\rho(x,h)=(Nx,h-m)$ for all $x\in\mathbb{R}^{d-1}$ and $h\in\left[m-\frac{1}{2},m+\frac{1}{2}\right]$.
It remains to prove the theorem in the case $N=1$ and $m=0$. Write
$\sigma=\sigma_{3^{d-1}}\circ\sigma_{3^{d-1}-1}\circ\ldots\circ\sigma_{1}|_{\mathbb{Z}^{d-1}},\quad\text{where}\quad\sigma_{1},\sigma_{2},\ldots,\sigma_{3^{d-1}}\colon\frac{1}{2}\mathbb{Z}^{d-1}\to\frac{1}{2}\mathbb{Z}^{d-1}$
are given, together with corresponding vectors $p_{1},\ldots,p_{3^{d-1}}\in\left\{-1,0,1\right\}^{d-1}$, for the permutation $\sigma$ by Lemma 4. For each $k\in[3^{d-1}]$ let $\tau_{k}\colon\mathbb{R}^{d-1}\to\mathbb{R}^{d-1}$ be the affine bijection defined by
$\tau_{k}(x)=\frac{x}{2}+Tp_{k}-T\sum_{i=1}^{d-1}e_{i}\qquad\text{for all $x\in\mathbb{R}^{d-1}$.}$
For each $k\in[3^{d-1}]$ we let $\widetilde{\sigma}_{k}\colon\mathbb{Z}^{d}\to\mathbb{Z}^{d}$ be the permutation defined by $\widetilde{\sigma}_{k}(x)=\tau_{k}^{-1}\circ\sigma_{k}\circ\tau_{k}$. Note that property (i) of each $\sigma_{k}$ from Lemma 4 translates to saying that each $\widetilde{\sigma}_{k}$ restricts to a permutation of every discrete tile of the form
$\mathbb{Z}^{d-1}\cap\left(\left(-\frac{1}{2},\ldots,-\frac{1}{2}\right)+\left[0,6T\right]^{d-1}+6Tz\right)\quad\text{with}\quad z\in\mathbb{Z}^{d-1}.$
Thus, for each $k\in[3^{d-1}]$ the permutation $\widetilde{\sigma}_{k}\colon\mathbb{Z}^{d-1}\to\mathbb{Z}^{d-1}$ satisfies the conditions of statement ((*)6) inside Lemma 6 with $S=6T$. Lemma 6 then guarantees, for each $k\in[3^{d-1}]$, the existence of a $\beta_{d}(6T)$-bilipschitz mapping $\widetilde{\Upsilon}_{k}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ satisfying ((*)6) and ((*)6) from statement ((*)6) with $\sigma=\widetilde{\sigma}_{k}$. Defining, for each $k\in[3^{d-1}]$, a bijection $\widehat{\tau}_{k}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ by
$\widehat{\tau}_{k}(x,h)=\left(\tau_{k}(x),\frac{1}{2}h\right)\qquad\text{for all $x\in\mathbb{R}^{d-1}$ and $h\in\mathbb{R}$,}$
the desired mapping $\Upsilon$ may be taken as $\Upsilon=\Upsilon_{3^{d-1}}\circ\Upsilon_{3^{d-1}-1}\circ\ldots\circ\Upsilon_{1}$, where for each $k\in[3^{d-1}]$, we take $\Upsilon_{k}:=\widehat{\tau}_{k}\circ\widetilde{\Upsilon}_{k}\circ{\widehat{\tau}_{k}}^{-1}|_{\mathbb{R}^{d-1}\times\left[-\frac{1}{2},\frac{1}{2}\right]}$, so that $\operatorname{bilip}\left(\Upsilon_{k}\right)\leq\beta_{d}(6T)$. Note that for each $k\in[3^{d-1}]$ and $x\in\frac{1}{2}\mathbb{Z}^{d-1}$ we have
$\Upsilon_{k}(x,0)=\widehat{\tau}_{k}\left(\widetilde{\Upsilon}_{k}(\tau^{-1}_{k}(x),0)\right)=\widehat{\tau}_{k}\left(\widetilde{\sigma}_{k}\circ\tau^{-1}_{k}(x),0\right)=\left(\sigma_{k}(x),0\right).\qed$
6 Threading a bilipschitz cord through the channel.
In the present section we extend a given bilipschitz mapping of two ‘horizontal’ hyperplanes in $\mathbb{R}^{d}$ and all integer lattice points between them, under the additional conditions that the mapping doesn’t break obvious topological requirements for extendability and coincides with the identity on the two hyperplanes. The section title comes from the early stages of the development of this work, where we took the dimension $d=2$ and the two horizontal hyperplanes, or lines, to have heights ±$\pm\frac{1}{2}$. In this special case there is only ‘one layer’, ×$\mathbb{Z}\times\left\{0\right\}$, of integer lattice points between the two lines and we imagined our task of extending $f$ to be like threading a bilipschitz cord through the points $f(x,0)$ for $x\in\mathbb{Z}$ in the correct order to determine the extension on ×$\mathbb{R}\times\left\{0\right\}$.
[…]
The reader may wish to consult the sketch proof of Theorem 1 in the introduction before proceeding.
Lemma 7.
Let $d,H\in\mathbb{N}$, $d\geq 2$, $0<s\leq\frac{1}{4}$, $N\geq(2\sqrt{d}/s)^{3}H^{\frac{1}{d-1}}$,
$X_{1},X_{2},\ldots,X_{H}\subseteq\mathbb{R}^{d-1}\times\left[\frac{1}{2}+s,H+\frac{1}{2}-s\right]$
be pairwise disjoint and suppose that $X:=\displaystyle\bigcup_{m=1}^{H}X_{m}$ is $s$-separated. Then there exists a $2^{8}N^{2}H^{2}$-bilipschitz mapping $\Psi\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that
  1. (i)
    $\Psi(x)=x$ for all $x\in\mathbb{R}^{d-1}\times\left(\mathbb{R}\setminus\left(\frac{1}{2},H+\frac{1}{2}\right)\right)$.
  2. (ii)
    $\Psi(X_{m})\subseteq\left(\frac{1}{N}\mathbb{Z}^{d-1}\setminus\mathbb{Z}^{d-1}\right)\times\left\{m\right\}$ for $m=1,2,\ldots,H$.
  3. (iii)
    $\left\|\operatorname{proj}_{\mathbb{R}^{d-1}}\circ\Psi(x)-\operatorname{proj}_{\mathbb{R}^{d-1}}(x)\right\|\leq s^{2}$ for all $x\in X$.
Proof.
[…]
Construction of $\Xi$.
We will define a 161616-bilipschitz mapping $\Xi\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ with the following properties
  1. (A)
    $\Xi(x)=x$ for all $x\in\mathbb{R}^{d-1}\times\left(\mathbb{R}\setminus\left(\frac{1}{2},H+\frac{1}{2}\right)\right)$,
  2. (B)
    $\left\|\Xi(x)-x\right\|\leq s^{2}$ for all $x\in X$.
  3. (C)
    $\operatorname{proj}_{\mathbb{R}^{d-1}}\circ\Xi(X)\subseteq\frac{1}{N}\mathbb{Z}^{d-1}\setminus\mathbb{Z}^{d-1}$ and $\operatorname{proj}_{\mathbb{R}^{d-1}}\circ\Xi|_{X}$ is injective.
  4. (D)
    $\operatorname{proj}_{d}\circ\Xi(x)=\operatorname{proj}_{d}(x)$ for all $x\in X$.
Let $Q:=\left[0,\frac{s^{2}}{\sqrt{d-1}}\right)^{d-1}$ and for each $x\in X$ let $\Phi(x)\in\mathbb{Z}^{d-1}$ be defined by the condition
$\operatorname{proj}_{\mathbb{R}^{d-1}}(x)\in Q+\frac{s^{2}}{\sqrt{d-1}}\Phi(x).$
For each $z\in\mathbb{Z}^{d-1}$ we have that $\bigcup_{x\in\Phi^{-1}(\left\{z\right\})}B(x,s/2)$ is a pairwise disjoint union of balls all contained in the set $\left(\overline{B}_{\mathbb{R}^{d-1}}(Q,s/2)+\frac{s^{2}z}{\sqrt{d-1}}\right)\times\left[\frac{1}{2},H+\frac{1}{2}\right]$. Hence, by comparing volumes and writing $\alpha_{d}$ for the $d$-dimensional volume of the unit ball in $\mathbb{R}^{d}$, we get
$\left|\Phi^{-1}(\left\{z\right\})\right|\leq\frac{(s+s^{2})^{d-1}H}{\alpha_{d}\left(s/2\right)^{d}}\leq\frac{(2s)^{d-1}2^{d}H}{\alpha_{d}s^{d}}\leq\frac{2^{d-1}\sqrt{d^{d}}H}{s}.$ (7)
On the other hand we have that
$\left|\left(Q+\frac{s^{2}}{\sqrt{d-1}}z\right)\cap\left(\frac{1}{N}\mathbb{Z}^{d-1}\setminus\mathbb{Z}^{d-1}\right)\right|\geq\left(\frac{s^{2}/\sqrt{d}}{1/N}-1\right)^{d-1}-1\geq\frac{1}{2}\left(\frac{Ns^{2}}{2\sqrt{d}}\right)^{d-1}.$ (8)
The lower bound on $N$ in the hypothesis is used for the last inequality. It also ensures that the final quantity of (8) is at least that of (7). Therefore, there exists for each $z\in\mathbb{Z}^{d-1}$ an injective mapping
$\iota_{z}\colon\Phi^{-1}(\left\{z\right\})\to\left(Q+\frac{s^{2}}{\sqrt{d-1}}z\right)\cap\left(\frac{1}{N}\mathbb{Z}^{d-1}\setminus\mathbb{Z}^{d-1}\right).$
For each $x\in X$ and a radius $r_{x}=r_{x}(s)$ to be determined later, we let
$y_{x}:=(\iota_{\Phi(x)}(x),x_{d})\in\mathbb{R}^{d},\qquad\mathfrak{U}_{x}:=B([x,y_{x}],r_{x}).$
Observe that for each $x\in X$ the line segment $[x,y_{x}]$ is parallel to the hyperplane $\mathbb{R}^{d-1}\times\left\{0\right\}$ and has length
$\left\|y_{x}-x\right\|\leq\operatorname{diam}\left(Q\right)=s^{2}.$ (9)
Therefore whenever $x,x^{\prime}\in X$ and $\operatorname{proj}_{\mathbb{R}^{d-1}}\mathfrak{U}_{x}\cap\operatorname{proj}_{\mathbb{R}^{d-1}}{\mathfrak{U}_{x^{\prime}}}\neq\emptyset$ we have
$\left\|\operatorname{proj}_{\mathbb{R}^{d-1}}(x^{\prime})-\operatorname{proj}_{\mathbb{R}^{d-1}}(x)\right\|\leq 2s^{2}+r_{x}+r_{x^{\prime}}\leq 3s^{2},$
where the last inequality comes from imposing the condition $r_{x}(s)\leq s^{2}/2$ for every $x\in X$. Therefore, whenever $x$ and $x^{\prime}$ are distinct points of the $s$-separated set $X$ and $\operatorname{proj}_{\mathbb{R}^{d-1}}\mathfrak{U}_{x}\cap\operatorname{proj}_{\mathbb{R}^{d-1}}{\mathfrak{U}_{x^{\prime}}}\neq\emptyset$ we have
$\left|\operatorname{proj}_{d}(x^{\prime})-\operatorname{proj}_{d}(x)\right|\geq s-3s^{2}\geq\frac{s}{4}\geq s^{2}\geq 2\sup_{w\in X}r_{w}(s).$
We deduce that the collection $\left(\mathfrak{U}_{x}\right)_{x\in X}$ is pairwise disjoint. Moreover, the condition $X\subseteq\mathbb{R}^{d-1}\times\left[\frac{1}{2}+s,H+\frac{1}{2}-s\right]$ and the fact that each $[x,y_{x}]$ is parallel to $\mathbb{R}^{d-1}\times\left\{0\right\}$ ensure that each set $\mathfrak{U}_{x}$ in this collection is contained in $\mathbb{R}^{d-1}\times\left(\frac{1}{2},H+\frac{1}{2}\right)$.
Let $\Xi\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ be the 161616-bilipschitz mapping given by the application of Theorem 3.1 to $X$, $(y_{x})_{x\in X}$ and $(r_{x})_{x\in X}$. Here we need one last condition on $r_{x}(s)$, namely $r_{x}(s)\leq\left\|y_{x}-x\right\|/2$ for all $x\in X$ with $y_{x}\neq x$; the choice
$r_{x}(s):=\begin{cases}\frac{1}{2}\left\|y_{x}-x\right\|&\text{if $y_{x}\neq x$,}\\ \frac{1}{2}s^{2}&\text{otherwise}\end{cases}$
satisfies both conditions we required. The properties (A)–(D) of $\Xi$ are now easily verified.
Construction of $\Omega$.
We construct a $16N^{2}H^{2}$-bilipschitz mapping $\Omega$ with the following properties:
  1. (I)
    $\Omega(x)=x$ for all $x\in\mathbb{R}^{d-1}\times\left(\mathbb{R}\setminus\left(\frac{1}{2},H+\frac{1}{2}\right)\right)$.
  2. (II)
    $\operatorname{proj}_{\mathbb{R}^{d-1}}\circ\Omega\circ\Xi(x)=\operatorname{proj}_{\mathbb{R}^{d-1}}\circ\Xi(x)$ for all $x\in X$.
  3. (III)
    $\operatorname{proj}_{d}\circ\Omega\circ\Xi(x)=m$ for all $x\in X_{m}$ and $m=1,2,\ldots,H$.
For each $m\in[H]$ and $x\in X_{m}$ let
$v_{x}$ $:=\Xi(x),$ $y_{x}:=(\operatorname{proj}_{\mathbb{R}^{d-1}}\circ\Xi(x),m)\in\mathbb{R}^{d},$
$r_{x}$ :=$:=\begin{cases}\min\left\{\frac{1}{2N},\frac{\left\|y_{x}-v_{x}\right\|}{2}\right\}&\text{if }v_{x}\neq y_{x},\\ \frac{1}{2N}&\text{if }v_{x}=y_{x},\end{cases}$ $\mathfrak{U}_{x}:=B([v_{x},y_{x}],r_{x}).$
The choice of these objects, together with property (C) of $\Xi$, ensures that the collection $(\mathfrak{U}_{x})_{x\in X}$ is pairwise disjoint. Moreover, the hypotheses on $X$, the bounds $r_{x}<1/N<s$ and (D) ensure that each set $\mathfrak{U}_{x}$ with $x\in X$ is contained in $\mathbb{R}^{d-1}\times\left(\frac{1}{2},H+\frac{1}{2}\right)$.
Let the $16N^{2}H^{2}$-bilipschitz mapping $\Omega\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ be given by Theorem 3.1 applied to $\left\{v_{x}\colon x\in X\right\}$ in place of $X$, $v_{x}$ in the role of $x$, $(y_{x})_{x\in X}$ and $(r_{x})_{x\in X}$. The properties (I)–(III) of $\Omega$ are clear from the construction.
Properties of $\Psi:=\Omega\circ\Xi$.
Setting $\Psi:=\Omega\circ\Xi$ we get that $\Psi$ is $2^{8}N^{2}H^{2}$-bilipschitz. The properties (i)–(iii) of $\Psi$ follow easily from (A)–(D) for $\Xi$ and (I)–(III) for $\Omega$. ∎ We are now ready to prove Theorem 1:
Proof of Theorem 1.
Let the
$2^{8}N^{2}H^{2}\text{-bilipschitz mapping }\Psi\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$
be given by the conclusion of Lemma 7 applied to
$s:=\frac{1}{4L},\quad X_{m}:=f\left(\mathbb{Z}^{d-1}\times\left\{m\right\}\right),\quad X:=f(\mathbb{Z}^{d-1}\times[H]),\quad N:=\lfloor 2(2\sqrt{d}(4L))^{3}H^{\frac{1}{d-1}}\rfloor\in\mathbb{N}.$ (10)
Let $m\in[H]$. Then we define
$\xi_{m}\colon\mathbb{Z}^{d-1}\to\frac{1}{N}\mathbb{Z}^{d-1}\setminus\mathbb{Z}^{d-1}$
by
$\xi_{m}(x)=\operatorname{proj}_{\mathbb{R}^{d-1}}\circ\Psi\circ f(x,m),\qquad x\in\mathbb{Z}^{d-1}.$
Using properties (i) and (ii) of $\Psi$ from Lemma 7, we have
$\left\|\xi_{m}(x)-x\right\|=\left\|\Psi\circ f(x,m)-(x,m)\right\|\\ \leq\left\|\Psi\circ f(x,m)-\Psi\circ f\biggl{(}x,\frac{1}{2}\biggr{)}\right\|+\left\|\biggl{(}x,\frac{1}{2}\biggr{)}-(x,m)\right\|\\ \leq 2^{8}LN^{2}H^{3}+H\leq\lfloor 2^{9}LN^{2}H^{3}\rfloor=:T,\qquad\text{for all $x\in\mathbb{Z}^{d-1}$.}$ (11)
Since $N\in\mathbb{N}$ we have that $\mathbb{Z}^{d-1}\subseteq\frac{1}{N}\mathbb{Z}^{d-1}$, so, in particular, $\xi_{m}$ is defined on a subset of $\frac{1}{N}\mathbb{Z}^{d-1}$. We now extend $\xi_{m}$ to a permutation $\sigma_{m}\colon\frac{1}{N}\mathbb{Z}^{d-1}\to\frac{1}{N}\mathbb{Z}^{d-1}$ as follows: Since each $\xi_{m}$ is injective we may consider its inverse $\xi_{m}^{-1}\colon\xi_{m}(\mathbb{Z}^{d-1})\to\mathbb{Z}^{d-1}$. Remembering that $\mathbb{Z}^{d-1}\cap\xi_{m}(\mathbb{Z}^{d-1})=\emptyset$, we may define $\sigma_{m}\colon\frac{1}{N}\mathbb{Z}^{d-1}\to\frac{1}{N}\mathbb{Z}^{d-1}$ by
$\sigma_{m}(x)=\begin{cases}\xi_{m}(x)&\text{if }x\in\mathbb{Z}^{d-1},\\ \xi^{-1}_{m}(x)&\text{if }x\in\xi_{m}(\mathbb{Z}^{d-1}),\\ x&\text{if }x\in\frac{1}{N}\mathbb{Z}^{d-1}\setminus\left(\mathbb{Z}^{d-1}\cup\xi_{m}\left(\mathbb{Z}^{d-1}\right)\right).\end{cases}$
It is clear that $\sigma_{m}\colon\frac{1}{N}\mathbb{Z}^{d-1}\to\frac{1}{N}\mathbb{Z}^{d-1}$ is a permutation extending $\xi_{m}$, $\sigma_{m}\circ\sigma_{m}=\operatorname{id}$ and that $\left\|\sigma_{m}(x)-x\right\|\leq T$, defined by (11), for all $x\in\frac{1}{N}\mathbb{Z}^{d-1}$. Let $\Upsilon\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ be defined by
$\Upsilon(x)=\begin{cases}\Upsilon_{m}(x)&\text{if }x\in\mathbb{R}^{d-1}\times\left[m-\frac{1}{2},m+\frac{1}{2}\right],\,m\in[H],\\ x&\text{if }x\in\mathbb{R}^{d-1}\times\left(\mathbb{R}\setminus\left(\frac{1}{2},H+\frac{1}{2}\right)\right),\end{cases}$
where for each $m\in[H]$ the $N\beta_{d}\left(6NT\right)^{3^{d-1}}$- bilipschitz mapping
$\Upsilon_{m}\colon\mathbb{R}^{d-1}\times\left[m-\frac{1}{2},m+\frac{1}{2}\right]\to\mathbb{R}^{d-1}\times\left[m-\frac{1}{2},m+\frac{1}{2}\right]$
is given by Theorem 5.1 for the permutation $\sigma_{m}\colon\frac{1}{N}\mathbb{Z}^{d-1}\to\frac{1}{N}\mathbb{Z}^{d-1}$. By Lemma 1 applied to $\Upsilon$ and the collection of open sets $\left(\mathbb{R}^{d-1}\times\left(m-\frac{1}{2},m+\frac{1}{2}\right)\right)_{m\in[H]}$ together with $\mathbb{R}^{d-1}\times\left(-\infty,\frac{1}{2}\right)$ and $\mathbb{R}^{d-1}\times\left(H+\frac{1}{2},\infty\right)$ we verify that $\Upsilon$ is bilipschitz with
$\operatorname{bilip}\left(\Upsilon\right)\leq N\beta_{d}\left(6NT\right)^{3^{d-1}}.$
Further, for all $x\in\mathbb{Z}^{d-1}$ and $m\in[H]$, we have
$\Upsilon\circ\Psi\circ f(x,m)=\Upsilon(\xi_{m}(x),m)=(\sigma_{m}\circ\xi_{m}(x),m)=(x,m).$ (12)
The desired extension $F\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ may now be defined as
$F:=\Psi^{-1}\circ\Upsilon^{-1}.$
Applying this mapping to both sides of (12) we verify that $F(x,m)=f(x,m)$ for all $x\in\mathbb{Z}^{d-1}$ and $m\in[H]$. Moreover, it is clear that $F$ coincides with the identity on $\mathbb{R}^{d-1}\times\left\{\frac{1}{2},H+\frac{1}{2}\right\}$. Thus, $F$ is an extension of $f$. Finally, note that
$\operatorname{bilip}(F)\leq\operatorname{bilip}(\Psi)\cdot\operatorname{bilip}(\Upsilon)\leq 2^{38}d^{9/2}L^{9}H^{2+\frac{3}{d-1}}\cdot\beta_{d}\left(2^{42}d^{9/2}L^{10}H^{3+\frac{3}{d-1}}\right)^{3^{d-1}},$
where in the last inequality we exploit the monotone increasing property of $\beta_{d}$ from Lemma 6 and apply the bounds $N\leq 2^{10}d^{3/2}L^{3}H^{\frac{1}{d-1}}$ and $T\leq 2^{9}LN^{2}H^{3}$ from (10) and (11). ∎